# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v2

@@ 2.2. Relation to the periodic polymer
For $\mathbf{Y}=(\mathbf{Y}_{i})_{i\in\mathbb{Z}},\mathbf{Y}^{\prime}=(\mathbf{Y}^{\prime}_{i})_{i\in\mathbb{Z}}\in(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ satisfying $\mathfrak{s}(\mathbf{Y}_{1})=\mathfrak{s}(\mathbf{Y}^{\prime}_{1})$, $\mathfrak{s}(\mathbf{Y}_{2})=\mathfrak{s}(\mathbf{Y}_{2}^{\prime})$, we have that $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})$ if and only if $\mathbf{Y}_{1}=\mathbf{Y}^{\prime}_{1},\mathbf{Y}_{2}=\mathbf{Y}_{2}^{\prime}$.
Before proving Lemma 3.1, we first use Lemmas A.8 and 3.1 to complete the proof of Theorem 1.6:
Proof of Theorem 1.6..
Using the partition function invariance under the operator $\mathcal{P}_{i+1}\mathcal{P}_{i}\mathcal{P}_{i+1}$ yields:
$H\bigl{(}\mathbf{W}_{k}\bigr{)}H\bigl{(}\mathbf{W}_{k+1}\bigr{)}H\bigl{(}\mathbf{W}_{k+2}\bigr{)}=$
[…]
$\qquad H\Bigl{(}D\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\bigr{)}\Bigr{)}.$ (3.4)
On the other hand, we can also apply the partition function invariance under the operator $\mathcal{P}_{i}\mathcal{P}_{i+1}\mathcal{P}_{i}$:
$H\bigl{(}\mathbf{W}_{k}\bigr{)}H\bigl{(}\mathbf{W}_{k+1}\bigr{)}H\bigl{(}\mathbf{W}_{k+2}\bigr{)}=$
[…]
$\qquad H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k}))\Bigr{)}.$ (3.5)
$H\Bigl{(}T\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}H\Bigl{(}T\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\Bigr{)}$
$=H\Bigl{(}T\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}.$
By repeated application of Lemma A.4, we have that
$\mathfrak{s}(T(T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}))=\mathfrak{s}(\mathbf{W}_{k+2})$
$=\mathfrak{s}(T(T(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})),T(\mathbf{W}_{k+1},\mathbf{W}_{k}))),\quad\text{and}$
$\mathfrak{s}(T(D(T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k})),D(\mathbf{W}_{k+2},\mathbf{W}_{k+1})))$
$=\mathfrak{s}(\mathbf{W}_{k+1})=\mathfrak{s}(T\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)})).$
Proof of Lemma 3.1..
Immediately, the backwards direction is clear. Then, it remains to show that, under the assumption that $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})\implies\mathbf{Y}_{1}=\mathbf{Y}^{\prime}_{1},\mathbf{Y}_{2}=\mathbf{Y}_{2}^{\prime}$. Let us prove by backwards induction on $k=N,\ldots,1$ that $Y_{1,k}=Y^{\prime}_{1,k}$ and $Y_{2,k}=Y^{\prime}_{2,k}$.
As a base case, let’s show this for $k=N$. From Lemma 2.6, note that
$\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N}$ $=Z^{\mathbf{Y}}(2,N\mid 1,1)$
$=e^{Y_{2,N}}\left(Z^{\mathbf{Y}}(1,N\mid 1,1)+Z^{\mathbf{Y}}(2,N-1\mid 1,1)\right)$
$=e^{Y_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y_{1,i}}+\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N-1}\Biggr{)}.$ (3.6)
Similarly, we conclude that:
$\left(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}^{\prime}_{2})\right)_{1,N}$ $=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,1)$
$=e^{Y^{\prime}_{2,N}}\left(Z^{\mathbf{Y}^{\prime}}(1,N\mid 1,1)+Z^{\mathbf{Y}^{\prime}}(2,N-1\mid 1,1)\right)$
$=e^{Y^{\prime}_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y^{\prime}_{1,i}}+\left(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}^{\prime}_{2})\right)_{1,N-1}\Biggr{)}.$ (3.7)
By assumption, we have $H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})=H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime})$ and $\prod_{i=1}^{N}e^{Y_{1,i}}=e^{\mathfrak{s}(Y_{1})}=e^{\mathfrak{s}(Y^{\prime}_{1})}=\prod_{i=1}^{N}e^{Y^{\prime}_{1,i}}$. Equating (3.6) and (3.7) and canceling the nonzero expressions in parenthesis yields $Y_{2,N}=Y^{\prime}_{2,N}$. Next, we see that $Y^{\prime}_{1,N}=Y_{1,N}$, since:
$(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2}))_{(N,N)}=Z^{\mathbf{Y}}(2,N\mid 1,N)=e^{Y_{1,N}+Y_{2,N}},$ (3.8)
and
$(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime}))_{(N,N)}=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,N)=e^{Y^{\prime}_{1,N}+Y^{\prime}_{2,N}}.$ (3.9)
To finish the proof, simply notice that the proof of the above only required $\prod_{i=1}^{N}e^{Y^{\prime}_{i}}=\prod_{i=1}^{N}e^{Y_{i}}$. Alternatively, if we have $\prod_{i=1}^{k}e^{Y_{1,i}}=\prod_{i=1}^{k}e^{Y_{2,i}}$ for some $k\in\llbracket 1,N\rrbracket$, then the same proof shows that $Y_{1,k}=Y_{1,k}^{\prime}$ and $Y_{2,k}=Y_{2,k}^{\prime}$. This follows by backwards induction, since
$\prod_{i=1}^{k}e^{Y_{1,i}}=\mathfrak{s}(\mathbf{Y}_{1})\Bigl{(}\prod_{i=k+1}^{N}e^{Y_{1,i}}\Bigr{)}^{-1}.\qed$
4. Burke property and permutation invariance
In this section, we prove Theorems 1.7 and 1.8. We first need two intermediate lemmas. In the first, we generalize the Burke property in [CGS24, Proposition 5.12] to an inhomogeneous environment.
Proposition 4.1 (Log-inverse-gamma Burke property).
Let $\lambda>0$, and let $\mathbf{a},\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences such that $a_{i}+b_{j}>0$ for $i,j\in\mathbb{Z}_{N}$. Let $\mathbf{X}_{1},\mathbf{X}_{2}$ be independent sequences of independent random variables such that $X_{i,j}\sim(a_{i}+b_{j},\lambda)$ (see Definition 1.15). Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})).$
Proof.
Using Lemmas A.4-A.5 in the penultimate equality below, the joint density $\mathbb{P}(\mathbf{X}_{1}\in d\mathbf{x}_{1},\mathbf{X}_{2}\in d\mathbf{x}_{2})$ is proportional to
$=\exp\Biggl{(}-a_{1}\sum_{i\in\mathbb{Z}_{N}}x_{1,i}-a_{2}\sum_{i\in\mathbb{Z}_{N}}x_{2,i}-\sum_{i\in\mathbb{Z}_{N}}b_{i}(x_{1,i}+x_{2,i})-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{1,i}}+e^{-x_{2,i+1}})\Biggr{)}$
$=\exp\Biggl{(}-a_{1}\sum_{i\in\mathbb{Z}_{N}}T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-a_{2}\sum_{i\in\mathbb{Z}_{N}}D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-\sum_{i\in\mathbb{Z}_{N}}b_{i}(T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+D_{i}(\mathbf{x}_{1},\mathbf{x}_{2}))$
$=\exp\Biggl{(}-\sum_{i\in\mathbb{Z}_{N}}\Bigl{(}(a_{1}+b_{i})T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+(a_{2}+b_{i})D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})\Bigr{)}-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr{)},$
Thus, the density is preserved under our transformation, so all we must show is that the determinant of the Jacobian matrix of the transformation has absolute value 1. This follows from Lemma A.7. ∎
We next prove a lemma that connects the condition stated in Theorem 1.7 to the condition in Theorem 1.4.
Lemma 4.2.
@@ 4. Burke property and permutation invariance
Let $\sigma:\mathbb{Z}\to\mathbb{Z}$ be a finite permutation, decomposed as a product of nearest-neighbor transpositions $(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$, where $k_{1},\ldots,k_{n}\in\mathbb{Z}$. Assume that this decomposition has the minimal number of terms. Then, if $x\in\mathbb{Z}$ satisfies $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, then $x\neq k_{\ell}+1$ for all $1\leq\ell\leq n$. Similarly, if $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$, then $x\neq k_{\ell}$ for all $1\leq\ell\leq m$.
@@
For an arbitrary finite permutation $\sigma$, decompose it as a product of nearest-neighbor transpositions $\sigma=(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$. Take this decomposition to have the minimal number of terms. Then, we consider the action of the map $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}$ on the weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}$, as in Theorem 1.4. By Lemma 4.2, the condition (1.11) and the minimality of the decomposition ensures that the conditions of Theorem 1.4 are met, so
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U)=Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U),\quad\text{for all }(U,V)\in\mathcal{U}^{\sigma}.$
Furthermore, by successively applying the Burke property in Proposition 4.1, we see that each operator $\mathcal{P}_{k}$ swaps the column parameters in columns $k$ and $k+1$, so we have that $\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}^{\mathbf{a},\mathbf{b}}\overset{d}{=}\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}$. This completes the proof. ∎
@@
This is proved similarly as Theorem 1.7. Consider the independent weights in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and extend these periodically. Consider first endpoint pairs $(U,V)$ that consist of points lying in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and apply the partition function invariance in Theorem 1.4 and the Burke property in Proposition 4.1 (with Lemma 4.2 to connect the two conditions) to both rows and columns, with $N=2M+1$. The result follows by sending $M\to\infty$. ∎
Appendix A Inputs from previous work and technical inputs
A.1. Basic algebraic properties
In the present work, we use a different formulation of the maps $D$ and $T$ than those introduced in [CGS24]; there, the relevant mappings were defined as
$\mathfrak{T}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{1,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i,j]}-X_{1,[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i+1,j]-X_{1,[i+1,j]}}}}\Biggr{)},\quad\text{and}$
$\mathfrak{D}_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{2,i}+\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i,j]}-X_{1,(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i-1,j]}-X_{1,(i-1,j]}}}\Biggr{)}.$
In [CGS24], these maps were denoted as $T^{N,2}$ and $D^{N,2}$, respectively. It is important to note that this is equivalent to our definition of the maps defined in (1.2) up to a shift:
$D(\mathbf{X}_{1},\mathbf{X}_{2})=\mathfrak{D}(\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}),\quad\text{and}\quad T(\mathbf{X}_{1},\mathbf{X}_{2})=\tau_{-1}\mathfrak{T}(\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}).$ (A.1)
where, for $k\in\mathbb{Z}_{N}$, $\tau_{k}:\mathbb{R}^{\mathbb{Z}_{N}}\rightarrow\mathbb{R}^{\mathbb{Z}_{N}}$ is the shift map: $(\tau_{k}X)_{i}=X_{i+k}$ for all $i\in\mathbb{Z}_{N}$.

### v3

@@ 3. Invariance of polymer partition functions
Let $\sigma:\mathbb{Z}\to\mathbb{Z}$ be a finite permutation, decomposed as a product of nearest-neighbor transpositions $(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)$, where $k_{1},\ldots,k_{m}\in\mathbb{Z}$. Assume that this decomposition has the minimal number of terms. Then, if $x\in\mathbb{Z}$ satisfies $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, then $x\neq k_{\ell}+1$ for all $1\leq\ell\leq m$. Similarly, if $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$, then $x\neq k_{\ell}$ for all $1\leq\ell\leq m$.
@@
Let $\sigma\in\mathcal{S}_{\mathbb{Z}}$, and let $(U,V)\in\mathcal{W}^{\sigma}$ (recalling the definition of this set in (1.10)). Denote the points of $U$ and $V$ as $U=\bigl((a_{i},b_{i})\bigr)_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl((c_{i},d_{i})\bigr)_{i\in\llbracket 1,k\rrbracket}$. Let $\sigma\in\mathcal{S}_{\mathbb{Z}}$. By assumption that $(U,V)\in\mathcal{W}^{\sigma}$, Lemma 3.3 implies that the conditions of Proposition 3.2 are satisfied so that, for each $i\in\llbracket 1,k\rrbracket$,
$Z^{\mathbf{X}}(c_{j},d_{j}\mid a_{i},b_{i})=Z^{\mathcal{P}_{\sigma}\mathbf{X}}(c_{j},d_{j}\mid a_{i},b_{i}).$ (3.1)
To extend this to the multi-path result, we use the Lindström-Gessel-Viennot (LGV) lemma (cf. [AZ18, Chapter 32], originally from [LIN73, GV85]) to prove the theorem. Because LGV applies only to edge-weighted graphs, we need to transfer our vertex weights to edge weights in the following way: for each vertex $(r,s)$, split it into $(r,s)_{\text{in}}$ and $(r,s)_{\text{out}}$, with edge $\bigl((r,s)_{\text{in}},(r,s)_{\text{out}}\bigr)$ given the weight $\omega_{(r,s)}:=e^{X_{(r,s)}}$, and edges $\bigl((r,s)_{\text{out}},(r,s+1)_{\text{in}}\bigr)$ and $\bigl((r,s)_{\text{out}},(r+1,s)_{\text{in}}\bigr)$ given the weight 1. See Figure 5. Any path which started at $(r,s)$ starts at $(r,s)_{\text{in}}$, while any path ending at $(r,s)$ ends at $(r,s)_{\text{out}}$. Under this graph transformation, our partition function on the pre-transformed graph equals the sum over weight products along possible paths on the transformed diagram, so we can apply LGV.
[…]
Figure 5. Local transformation of a vertex-weighted node (left) with weight $X_{(r,s)}$ into an edge-weighted directed graph (right), where the internal edge carries weight $\omega_{(r,s)}=\exp(X_{(r,s)})$ and external edges carry weight 1. Paths in the original graph correspond to paths from “in” nodes to “out” nodes in the transformed graph, preserving path weights.
Define the matrix $M^{\mathbf{X}}:=\bigl(Z^{\mathbf{X}}(c_{j},d_{j}\mid a_{i},b_{i})\bigr)_{(i,j)\in\llbracket 1,m\rrbracket}$. By LGV, we have
$\det M^{\mathbf{X}}=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V_{\tau}}\ \forall\varphi\in\mathcal{S}_{m}}\operatorname{sgn}(\varphi)\prod\limits_{(i,j)\in P}e^{X_{(i,j)}}.$
Since $\mathcal{W}^{\sigma}\subseteq\Psi$, we have that, by definition of $\Psi$ (1.7), for $\varphi\neq\text{Id}$, $\Pi_{U,V_{\varphi}}=\varnothing$. Then, because $\operatorname{sgn}(\text{Id})=1$,
$\det M^{\mathbf{X}}=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V}}\prod\limits_{(i,j)\in P}e^{X_{(i,j)}}=Z^{\mathbf{X}}(V\mid U).$ (3.2)
By (3.1), for each $i,j$, we have $M_{i,j}^{\mathbf{X}}=M_{i,j}^{\mathcal{P}_{\sigma}\mathbf{X}}$. Therefore, by (3.2), we also have
$Z^{\mathcal{P}_{\sigma}\mathbf{X}}(V\mid U)=\det M^{\mathcal{P}_{\sigma}\mathbf{X}}=\det M^{\mathbf{X}}=Z^{\mathbf{X}}(V\mid U).\qed$
4. Burke property and permutation invariance
In this section, we prove Theorems 1.5 and 1.6. We first need two intermediate lemmas. In the first, we generalize the Burke property in [CGS26, Proposition 5.12] to an inhomogeneous environment.
Proposition 4.1 (Log-inverse-gamma Burke property).
Let $\lambda>0$, let $a_{1},a_{2}\in\mathbb{R}$, and let $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be a sequence such that $a_{m}+b_{i}>0$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$. Let $\mathbf{X}_{1},\mathbf{X}_{2}$ be independent sequences of independent random variables such that $X_{m,i}\sim\text{Log-Inv-Gamma}(a_{m}+b_{i},\lambda)$ (see Definition 1.13). Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(D(\mathbf{X}_{1},\mathbf{X}_{2}),T(\mathbf{X}_{1},\mathbf{X}_{2})).$
Proof.
Using Lemmas 2.1-2.2 in the penultimate equality below, the joint density $\mathbb{P}(\mathbf{X}_{1}\in d\mathbf{x}_{1},\mathbf{X}_{2}\in d\mathbf{x}_{2})$ is proportional to
$=\exp\Biggl(-a_{1}\sum_{i\in\mathbb{Z}_{N}}x_{1,i}-a_{2}\sum_{i\in\mathbb{Z}_{N}}x_{2,i}-\sum_{i\in\mathbb{Z}_{N}}b_{i}(x_{1,i}+x_{2,i})-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-x_{2,i}}+e^{-x_{1,i+1}})\Biggr)$
$=\exp\Biggl(-a_{1}\sum_{i\in\mathbb{Z}_{N}}D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-a_{2}\sum_{i\in\mathbb{Z}_{N}}T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})-\sum_{i\in\mathbb{Z}_{N}}b_{i}(T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+D_{i}(\mathbf{x}_{1},\mathbf{x}_{2}))$
$=\exp\Biggl(-\sum_{i\in\mathbb{Z}_{N}}\Bigl((a_{1}+b_{i})D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})+(a_{2}+b_{i})T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})\Bigr)-\lambda\sum_{i\in\mathbb{Z}_{N}}(e^{-D_{i}(\mathbf{x}_{1},\mathbf{x}_{2})}+e^{-T_{i}(\mathbf{x}_{1},\mathbf{x}_{2})})\Biggr).$
Thus, the density is preserved under our transformation, so all we must show is that the determinant of the Jacobian matrix of the transformation has absolute value 1. This was shown previously in [CGS26], recorded in the present paper as Lemma A.1. ∎
@@ 4. Burke property and permutation invariance
For an arbitrary finite permutation $\sigma$, Theorem 1.4 implies that
$Z^{\mathcal{P}_{\sigma}\mathbf{X}^{\mathbf{a},\mathbf{b}}}(V\mid U)=Z^{\mathbf{X}^{\mathbf{a},\mathbf{b}}}(V\mid U),\quad\text{for all }(U,V)\in\mathcal{W}^{\sigma}.$
Furthermore, by successively applying the Burke property in Proposition 4.1, we see that each operator $\mathcal{P}_{k}$ swaps the column parameters in columns $k$ and $k+1$, so we have that $\mathcal{P}_{\sigma}\mathbf{X}^{\mathbf{a},\mathbf{b}}\overset{d}{=}\mathbf{X}^{\sigma\mathbf{a},\mathbf{b}}$. This completes the proof. ∎
Proof of Theorem 1.6..
This is proved similarly as Theorem 1.5. Consider the independent weights in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and extend these periodically. Consider first end point pairs $(U,V)$ that consist of points lying in the finite grid $\llbracket-M,M\rrbracket\times\llbracket-M,M\rrbracket$, and apply the partition function invariance in Theorem 1.4 and the Burke property in Proposition 4.1 to both rows and columns, with $N=2M+1$. The result follows by sending $M\to\infty$. ∎
Appendix A Inputs from previous results and zero-temperature analogues
A.1. Additional algebraic properties
In [CGS26], index-wise shifted versions of our $T$ and $D$ maps were denoted by $T^{N,2}$ and $D^{N,2}$ respectively. To be precise, the definition of the mappings in that paper were
$T_{i}^{N,2}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{1,i}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i,j]}-X_{1,[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,[i+1,j]}-X_{1,[i+1,j]}}}\Biggr),\quad\text{and}$
$D_{i}^{N,2}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})$ $=X_{2,i}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i,j]}-X_{1,(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{X_{2,(i-1,j]}-X_{1,(i-1,j]}}}\Biggr).$
In the present paper, we subsumed the dependence on $N$. From the definitions given in the present paper (1.2), we can see that
$T(\mathbf{X}_{1},\mathbf{X}_{2})=\tau^{-1}T^{N,2}(\mathbf{X}_{2},\tau\mathbf{X}_{1}),\quad\text{and}\quad D(\mathbf{X}_{1},\mathbf{X}_{2})=D^{N,2}(\mathbf{X}_{2},\tau\mathbf{X}_{1}),$ (A.1)
where $\tau:\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}$ is the shift map defined by $(\tau\mathbf{X})_{i}=X_{i+1}$. With this in mind, we now cite two results from [CGS26], which follow from translating them into appropriate statements for the present paper via the connection in (A.1).